# Multi-Preference Lambda-weighted Listwise DPO for Small-Scale Model Alignment

### v5

@@ Abstract
Large language models (LLMs) demonstrate strong generalization across a wide range of language tasks, but often generate outputs that misalign with human preferences. Reinforcement Learning from Human Feedback (RLHF) addresses this by optimizing models toward human preferences using a learned reward function and reinforcement learning, yielding improved alignment but suffering from high computational cost and instability. Direct Preference Optimization (DPO) simplifies the process by treating alignment as a classification task over binary preference pairs, reducing training overhead while achieving competitive performance. However, it assumes fixed, single-dimensional preferences and only supports pairwise supervision. To address these limitations, we propose Multi-Preference Lambda-weighted Listwise DPO, which allows the model to learn from more detailed human feedback and flexibly balance multiple goals such as helpfulness, honesty, and fluency. Our method models full-ranked preference distributions rather than binary comparisons, enabling more informative learning signals. The $\lambda$ vector controls the relative importance of different alignment goals, allowing the model to generalize across diverse human objectives. During inference, $\lambda$ can be adjusted without retraining, providing controllable alignment behavior for downstream use. We also introduce a learned scheduler that dynamically samples performant $\lambda$ configurations to improve robustness. Notably, our method requires only 20GB of GPU memory for training, making it suitable for compute-constrained settings such as academic labs, educational tools, or on-device assistants. Experiments on 1B–2B scale models show that our method consistently outperforms standard DPO on alignment benchmarks while enabling efficient, controllable, and fine-grained adaptation suitable for real-world deployment.
Introduction
Large language models (LLMs) such as GPT (Brown et al. 2020) and LLaMA (Touvron and et al. 2023) have emerged as powerful backbones in natural language processing, demonstrating impressive capabilities across tasks such as question answering, summarization, and instruction following. However, despite their success, LLMs often produce outputs that do not align well with human expectations and values.
This misalignment frequently manifests as responses that are verbose, inaccurate, or unhelpful. It stems from their pretraining objective: next-token prediction, which emphasizes statistical co-occurrence over communicative intent, and inherently lacks exposure to explicit human feedback (Ouyang et al. 2022).
To address this, a line of alignment techniques has been developed to incorporate preference supervision. Among them, Reinforcement Learning from Human Feedback (RLHF) has become a standard paradigm (Christiano et al. 2017; Ouyang et al. 2022; Bai et al. 2022). By training a reward model from human comparisons and optimizing it via reinforcement learning (e.g., PPO (Schulman et al. 2017)), RLHF enables models to better match human values. However, the reliance on expensive reward modeling, unstable optimization dynamics, and high computational cost make RLHF difficult to scale and deploy reliably (Ziegler et al. 2019).
As a lightweight alternative, Direct Preference Optimization (DPO) has recently been proposed (Rafailov et al. 2023). DPO reframes alignment as a classification problem over pairwise preferences, directly optimizing model outputs to prefer preferred responses over less preferred ones. Without requiring an explicit reward model or reinforcement loop, DPO achieves strong alignment with reduced training overhead. Nevertheless, existing DPO methods typically assume a single, fixed alignment objective and binary supervision. This limits their applicability in practical scenarios, where alignment is inherently multi-dimensional—balancing objectives such as helpfulness, factuality, honesty, and fluency (Bai et al. 2022; Wu et al. 2021).
To accommodate this diversity, several controllable generation strategies have been explored, such as prompt tuning (Liu and et al. 2021), control tokens (Keskar et al. 2019), and multi-reward reinforcement learning (Lee et al. 2023). However, these approaches often require retraining under each new preference configuration, introduce architectural complexity, or generalize poorly in unseen alignment regimes.
In this work, we propose Multi-Preference Lambda-weighted Listwise DPO, a unified framework that extends the standard DPO formulation along two critical dimensions: supervision granularity and preference flexibility. Concretely, we generalize pairwise supervision to listwise preference modeling, enabling the model to learn from multiple ranked completions per prompt. This richer supervision signal allows the model to reason over global ranking structures, rather than relying on isolated binary comparisons. As a result, it improves gradient quality and better captures nuanced distinctions between high-quality and suboptimal responses. Simultaneously, we introduce a lambda-weighted composition mechanism, in which a preference vector $\lambda\in\Delta^{m}$ governs the relative importance of each alignment axis. This design allows a single model to flexibly interpolate between multiple objectives without retraining.
To further enhance adaptability, we develop a learned scheduling strategy that dynamically adjusts the sampling distribution over $\lambda$ vectors based on their empirical performance.
Empirical results on two strong 1B–2B scale LLMs (LLaMA and Qwen) across six alignment benchmarks demonstrate that our approach consistently outperforms standard DPO. Furthermore, by adjusting $\lambda$ during inference, our method supports personalized alignment without modifying model weights, making it highly practical for real-world deployment in education, conversational agents, and lightweight on-device applications.
Related Work
RLHF and DPO.
Reinforcement learning from human feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022) has become a widely used paradigm for aligning large language models (LLMs) with human preferences. It typically involves training a reward model using pairwise human comparisons and then optimizing the base language model using reinforcement learning algorithms such as Proximal Policy Optimization (PPO) (Schulman et al. 2017). This two-stage approach has demonstrated strong alignment performance, notably in systems like InstructGPT (Ouyang et al. 2022) and Claude (Bai et al. 2022). However, RLHF suffers from several limitations: reward models can be noisy or biased, PPO introduces high variance and instability during training, and the overall pipeline is computationally intensive and hard to scale (Ziegler et al. 2019; Christiano et al. 2017).
To reduce this complexity, direct preference optimization (DPO) (Rafailov et al. 2023) reframes preference alignment as a supervised classification problem. Instead of learning a reward model, DPO directly optimizes the policy to prefer preferred responses using a KL-regularized binary cross-entropy loss. This improves both training stability and implementation simplicity. Variants such as LLaMA-DPO (Liu et al. 2023) and Quark (Yao et al. 2023) have extended this framework with architectural tuning and reward quantization, showing strong performance in single-objective settings. However, most existing work in this area assumes fixed binary preferences and lacks the flexibility to accommodate multiple alignment goals or dynamically varying human objectives.
Listwise Preference Learning.
While pairwise preference modeling is common, listwise preference learning has been shown to better capture global ranking information (Burges et al. 2005; Xia et al. 2008). Instead of treating comparisons independently, listwise approaches model the entire ranking over a set of candidates, providing richer supervision and reducing sample complexity. In the context of language model alignment, UltraFeedback (Du et al. 2023) provides large-scale human feedback data with complete ranking annotations over multiple completions. This enables more expressive and fine-grained alignment signals beyond binary feedback.
Recent works such as GLOW (Xu et al. 2023) leverage such ranked data to train reward models using listwise objectives. These models are then used in RLHF-style pipelines to align policies. However, they still rely on intermediate reward learning and reinforcement optimization. In contrast, our method directly optimizes the policy using listwise cross-entropy loss, bypassing the need for explicit reward modeling. This not only simplifies the training pipeline but also provides more stable gradients and better captures the structure of human preferences.
Multi-Objective and Controllable Alignment.
Human alignment is inherently multi-faceted, involving trade-offs among objectives such as helpfulness, honesty, informativeness, and harmlessness. Prior work has explored multi-objective reinforcement learning frameworks (Korbak et al. 2023; Lee et al. 2023) that combine multiple reward signals through scalarization or decomposition. However, these methods typically require retraining or manual tuning of trade-off weights for each new configuration. Similarly, reward interpolation techniques (Wu et al. 2021) attempt to blend preferences using convex combinations of pretrained reward models, but they still involve separate optimization for each alignment setting.
To enable better generalization, our method integrates a weight vector $\lambda$ into the training process, allowing the model to internalize a continuous spectrum of preference trade-offs. This is conceptually related to universal value function approximators (Schaul et al. 2015), which generalize value functions across goals or reward vectors. In the alignment context, we treat $\lambda$ as a parameter controlling the target alignment behavior, enabling downstream control without model retraining.
Alternative strategies for controllable generation include prompt-based steering (Brown et al. 2020; Wei et al. 2022), control tokens (Keskar et al. 2019), and constrained decoding techniques (Lu et al. 2022). While effective in certain settings, these methods often struggle with consistency, lack semantic grounding in human-aligned dimensions, or require careful hand-engineering. Our approach differs in that it integrates multi-dimensional alignment directly into the learning objective, enabling fine-grained, interpretable control through $\lambda$ without modifying prompts, decoding strategies, or model architecture.
Preliminaries
We begin by reviewing the foundation of Direct Preference Optimization (DPO) and its connection to KL-regularized reinforcement learning from human feedback (RLHF). Our setting assumes the existence of a base language model $\pi_{\text{ref}}$, a supervised fine-tuned (SFT) model, and a collection of ranked or pairwise human preferences. We aim to train a target policy $\pi_{\theta}$ that better aligns with human preferences while remaining close to $\pi_{\text{ref}}$.
Pairwise Preference Modeling.
Let $x$ denote the prompt, and $y_{w}\succ y_{l}$ represent a human preference between a pair of completions: $y_{w}$ (the preferred answer) and $y_{l}$ (the less preferred one). These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$. A commonly adopted model is the Bradley-Terry (BT) model (Bradley and Terry 1952):
$p^{*}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r(x,y_{w}))}{\exp(r(x,y_{w}))+\exp(r(x,y_{l}))}$ (1)
Given a dataset of comparisons $\mathcal{D}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, we can learn the reward function $r(x,y)$ by minimizing the negative log-likelihood under the BT model:
$\mathcal{L}_{\text{BT}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma(r(x,y_{w})-r(x,y_{l}))\right]$ (2)
where $\sigma(z)=1/(1+\exp(-z))$ is the logistic sigmoid.
@@ Introduction
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r(x,y)-\beta\cdot\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right]$ (3)
where $\beta>0$ controls the strength of the KL constraint.
This objective encourages the policy to produce high-reward outputs while staying close to $\pi_{\text{ref}}$. However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO (Schulman et al. 2017) to optimize.
@@
$\beta(r(x,y_{w})-r(x,y_{l}))=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ (5)
@@
$-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (6)
@@
Figure 1: Overview of the $\lambda$-DPO architecture. The system integrates listwise preference modeling and a controllable $\lambda$ vector to align model outputs with multi-dimensional human preferences.
Multi-Preference Lambda-weighted Listwise DPO
Motivated by the limitations of pairwise DPO under a single fixed preference distribution, we propose two orthogonal enhancements: listwise supervision and simplex-weighted optimization. First, instead of relying on binary comparisons, we adopt listwise learning to utilize full human rankings over multiple outputs. This provides stronger supervision signals and captures richer preference structure. Second, we generalize DPO to support multi-objective alignment by introducing a simplex-weighted aggregation over multiple reward components. The resulting $\lambda$-DPO formulation enables flexible interpolation across alignment dimensions such as helpfulness, harmlessness, and fluency. Together, these extensions allow controllable and scalable alignment with more nuanced human feedback.
Listwise Preference Modeling.
@@
$\mathcal{L}_{\text{listwise}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$ (8)
Lambda-weighted Preference Composition.
@@
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$ (10)
@@
What does $\lambda$-DPO update do?

### v6

# Listwise Direct Preference Optimization with Multi-Dimensional Preference Mixing

@@ Abstract
Recent alignment methods based on Direct Preference Optimization (DPO) reformulate preference learning as supervised optimization over pairwise comparisons, offering improved efficiency and stability over reinforcement learning from human feedback (RLHF). However, existing DPO-style methods implicitly assume a single fixed preference objective, which limits their ability to model the structured and sometimes conflicting nature of real-world human judgments that span multiple preference dimensions. In this work, we propose Listwise Direct Preference Optimization ($\lambda$-DPO), a unified framework that simultaneously improves supervision granularity and preference flexibility. Instead of collapsing multi-dimensional preference signals into a single ranking, $\lambda$-DPO constructs a mixture of listwise preference distributions weighted by a preference vector $\lambda$ on the probability simplex, enabling a single model to internalize a continuous spectrum of preference trade-offs. To further improve robustness, we introduce a performance-driven stochastic $\lambda$ scheduler that adaptively samples preference weights based on empirical downstream performance, explicitly mitigating the risks of misspecification inherent to static weighting schemes. We evaluate our method across multiple model families and scales on six widely used benchmarks. Experimental results show the consistent improvement against baselines.
Listwise Direct Preference Optimization with Multi-Dimensional Preference Mixing
Yuhui Sun University of Alberta          Xiyao Wang University of Toronto          Zixi Li Zhejiang University
YiTian Ding School of Computer Science McGill University          Tianyang Ling University of Alberta          Jialuo Chen University of Toronto
Tianyi Yu Alibaba Group (Ant Group)          Zhenlong Yuan Chinese Academy of Sciences          Jinman Zhao University of Toronto jzhao@cs.toronto.edu
1 Introduction
Large language models (LLMs) have demonstrated strong generalization capabilities such as question instruction following Brown et al. (2020); Touvron et al. (2023); OpenAI (2024). Despite these advances, their behavior in real-world applications frequently remains misaligned with human expectations, manifesting in issues such as factual inaccuracies, unsafe responses, and failures to appropriately interpret user intent Bai et al. (2022); Ouyang et al. (2022); Lin et al. (2022). Unlike early formulations that treated alignment as the optimization of a single performance metric, practical alignment objectives are inherently multi-dimensional, requiring models to simultaneously balance factors such as helpfulness, factuality, safety, fluency, and task-specific constraints often with intrinsic tensions and trade-offs among them Christiano et al. (2017); Wu et al. (2021); Korbak et al. (2023). Such human preferences are difficult to compress into a single scalar reward or fixed objective function; instead, they are more naturally expressed through relative comparisons and rankings over multiple candidate responses Bradley and Terry (1952); Burges et al. (2005); Cui et al. (2023). Consequently, the central challenge of alignment lies not in learning a single static “optimal” answer, but in modeling and generalizing the structured manifestation of multi-dimensional human preferences over sets of candidate outputs.
Multiple methods have gradually emerged as a dominant paradigm for aligning LLMs with human expectations. Among them, reinforcement learning from human feedback (RLHF) optimizes generation policies using preference signals, typically via policy gradient methods such as Proximal Policy Optimization (PPO) Schulman et al. (2017) has demonstrated substantial improvements Ouyang et al. (2022). However, RLHF often suffers from high engineering and computational overhead, as well as instability arising from reward modeling errors and reinforcement learning dynamics. Direct Preference Optimization (DPO) Rafailov et al. (2023) reformulates alignment as a supervised learning problem, directly optimizing the policy using pairwise preference comparisons. Building upon the DPO framework, a growing body of work has proposed various extensions and variants such as CPO Xu et al. (2024), SimPO Meng et al. (2024), RRHF Yuan et al. (2023), ORPO Hong et al. (2024) and IPO Azar et al. (2024). These work modify the optimization objective, regularization strategy, or preference formulation to improve robustness and efficiency under different training regimes.
Existing preference-alignment methods typically rely on an implicit modeling assumption that human preferences induce a single, fixed ranking over candidate responses. Recently, several listwise extensions of DPO such as LiPO Liu et al. (2025), Soft Preference Optimization Sharifnassab et al. (2024) have been proposed to directly optimize ranking distributions over sets of candidate responses, primarily aiming to improve supervision density and training stability. However, these methods still fundamentally assume that, for a given prompt, candidate responses correspond to a semantically fixed ranking target. In practice, this assumption is typically implemented by collapsing multi-dimensional preference signals into a single ranking objective. Nevertheless, real-world human preference data do not always support this formulation. For example, in multi-dimensional preference datasets such as UltraFeedback, candidate responses to the same prompt are evaluated independently along multiple criteria which often yield conflicting judgments across dimensions Cui et al. (2023).
Based on the above observations, we propose multi-preference weighted listwise DPO ($\lambda$-DPO), a unified framework that simultaneously improves supervision granularity and preference flexibility. Specifically, instead of simplifying preference signals into binary pairwise comparisons, our method leverages the multi-candidate ranking information available for each prompt and directly learns to match a listwise preference distribution with the model-induced distribution. This formulation provides richer and lower-variance training signals while capturing global ranking relationships over candidate sets. To account for the inherently multi-dimensional nature of real-world alignment objectives, we further introduce a preference weight vector $\lambda$ defined on the probability simplex, which combines listwise preference distributions corresponding to different preference dimensions via a weighted mixture, yielding a controllable target distribution $p_{\lambda}$. This design allows a single model to internalize a continuous spectrum of preference trade-offs during training and enables inference-time controllable alignment through adjusting $\lambda$ without retraining. In addition, we design a learned $\lambda$ scheduling strategy that dynamically adapts the sampling distribution of $\lambda$ based on empirical performance, improving training robustness and generalization while remaining compatible with resource-constrained settings. In summary, this paper makes the following contributions:
  • We identify a fundamental limitation of existing DPO and its extensions, which rely on binary supervision and a single fixed alignment objective, making them insufficient to capture the structure of multi-dimensional human preferences. To address this, we propose a multi-preference weighted listwise DPO framework.
  • We empirically validate the effectiveness of $\lambda$-DPO across multiple model scales and evaluation benchmarks, demonstrating consistent improvements over existing DPO settings.
  • Through ablation and analysis experiments, we systematically study the effects of different preference weight configurations and the learned $\lambda$ scheduling strategy, shedding light on how preference weighting influences training stability and alignment performance.
Figure 1: Overall framework.
2 Related Work
Alignment and RLHF.
RLHF Ouyang et al. (2022) is a foundational paradigm for aligning LLMs. Its core approach learns a reward model from human preference comparisons and optimizes the generation policy using policy gradient methods such as REINFORCE Sutton et al. (1999), PPO Schulman et al. (2017) or REINFORCE++ Hu et al. (2025). While effective, RLHF typically incurs high computational cost and suffers from training instability. To address these limitations, recent work explores alternatives that bypass explicit value or reward modeling, such as GRPO Shao et al. (2024), or reformulate alignment as a supervised learning problem, such as DPO Rafailov et al. (2023).
DPO Variation.
DPO has inspired a growing family of alignment methods that refine its objective or optimization strategy. Representative variants include SimPO Meng et al. (2024) simplifies the DPO objective by removing explicit reference policies; ORPO Hong et al. (2024) unifies preference optimization with likelihood training; CTO Xu et al. (2024) and IPO Azar et al. (2024) reinterpret preference learning from contrastive or implicit policy optimization perspectives; StepDPO Lai et al. (2024); Xu et al. (2025) extends DPO to step-level supervision for multi-step reasoning; and RDPO Park et al. (2024) explicitly disentangles response length from quality to mitigate length bias in preference optimization. More recently, online DPO methods study in interactive or streaming settings, highlighting the crucial role of sampling strategies in stabilizing and improving online alignment performance Shi et al. (2025); Tu et al. (2025); Pang et al. (2024). Despite differences in formulation and optimization details, these methods largely share a common reliance on pairwise supervision and assume a single implicit alignment objective. Recent surveys Xiao et al. (2025) provide systematic overviews of DPO-style methods, highlighting both their efficiency advantages and shared modeling assumptions.
Listwise Preference Optimization.
Recent work has explored preference optimization from pairwise comparisons to listwise feedback, aiming to more fully exploit the relative information among multiple candidate responses for a given prompt. Representative approaches include LiPO Liu et al. (2025) and Soft Preference Optimization Sharifnassab et al. (2024) formulate alignment as directly matching model distributions to target ranking distributions over response lists. Ordinal Preference Optimization Zhao et al. (2025) optimizes differentiable ranking metrics, and listwise extensions of DPO that leverage soft rankings or best-of-$K$ feedback. Beyond language models, listwise preference optimization has also been applied in other domains, such as vision–language models Pesaran Zadeh et al. (2025) and diffusion-based generative models Bai et al. (2025).
3 Preliminaries
We briefly review preference modeling and DPO. These form the foundation of our approach.
Preference Modeling.
Let $x$ denote a prompt, and $y_{w}\succ y_{l}$ a human preference indicating that completion $y_{w}$ is preferred over $y_{l}$. We assume such comparisons are drawn from an underlying preference distribution $p^{*}(y_{w}\succ y_{l}\mid x)$. A common instantiation is the Bradley–Terry model Bradley and Terry (1952) that parameterizes preferences via a latent scoring function $r(x,y)$:
$p^{*}(y_{w}\succ y_{l}\mid x)$
$=$ $\frac{\exp(r(x,y_{w}))}{\exp(r(x,y_{w}))+\exp(r(x,y_{l}))}$
$=$ $\sigma\!\left(r(x,y_{w})-r(x,y_{l})\right)$ (1)
@@ 3 Preliminaries
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}$ $\big[r(x,y)$ (2)
$-\beta\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\big]$
where $\beta>0$ controls the strength of the KL constraint. This objective encourages the policy to produce high-reward outputs while staying close to $\pi_{\text{ref}}$. However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO Schulman et al. (2017) to optimize.
@@
$\beta(r(x,y_{w})-r(x,y_{l}))$
$=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ (4)
@@
$-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\bigg[\log\sigma$ $\bigg(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}$ (5)
$-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\bigg)\bigg]$
@@
4 Multi-Preference Lambda-weighted Listwise DPO
Motivated by the limitations of pairwise DPO under a single fixed preference distribution, we propose two orthogonal enhancements: listwise supervision and simplex-weighted optimization. First, instead of relying on binary comparisons, we adopt listwise learning to utilize full human rankings over multiple outputs. This provides stronger supervision signals and captures richer preference structure. Second, we generalize DPO to support multi-objective alignment by introducing a simplex-weighted aggregation over multiple reward components. The resulting $\lambda$-DPO formulation enables flexible interpolation across alignment dimensions such as helpfulness, harmlessness, and fluency. Together, these extensions allow controllable and scalable alignment with more nuanced human feedback. The overview pipeline is illustrated in Figure 1.
4.1 Listwise Preference Modeling
@@
$\mathcal{L}_{\text{listwise}}(\theta)=$ (7)
$-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$
4.2 Lambda-weighted Preference Composition
@@
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=$ (9)
$-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$
@@
4.3 What does $\lambda$-DPO update do?